Find the solution to the differential equation ysin(y) dx + x(sin y, ycos y) dy = 0

Detailed answer
Answer #1


o = Lр(4 sork -Aujү) х - хр (414x4 Q= Le(c-kN4 xp (Lk) н Leocht huis ah we Churraleshwe furth= (hikon No1,9) 2 x (Kimy-toy)+ M14,4)= y cosecy Nichiyla a cose2y lainy=y Lesy) Tregeneral solution Sanchiy) dy=c Y = constant Norcentaining of a J y cose